A target product has 8 simple inputs, 1 average input, and 9 complex inputs. There are 41 average outputs, 7 simple inquiries, 16 average master files, and 14 complex interfaces. Determine the unadjusted function points (UFP).
Answer #1

Answer:-

Simple Average Complex
Input 8 1 9
Output 41
Inquiry 7
Master Files 16
Interface 14

Therefore, Unadjusted function points = 8*3+1*4+9*6+41*5+7*3+16*10+14*10

                                                                   = 24+4+54+205+21+160+140 = 608

Hope it will helps. Please give Thumbs Up!! Thank you for posting the question, All the best.